# Atlas Gaussians Diffusion for 3D Generation with Infinite Number of Points

### v1

@@ 3.2 Stage 1: VAE
We design a VAE to link the low-dimensional latent space and the 3D space. The overall pipeline is shown in Figure 3.
Encoder. The encoder takes shape information as input and outputs a latent set $\boldsymbol{z}_{0}\in\mathbb{R}^{n\times d_{0}}$, which follows the latent set representation in [114]. Specifically, shape information is represented as a set of features $\mathcal{S}$. We use transformers with cross attention to extract the latent features, as $\boldsymbol{z}=\text{CrossAttn}(\bar{\boldsymbol{z}},\mathcal{S}),$ where $\bar{\boldsymbol{z}}$ is the initialization of the latent features. In this notation, We omit the feed-forward MLP of the transformers for simplicity. Intuitively, the query $\bar{\boldsymbol{z}}$ of the cross-attention is updated by aggregating information from key/value $\mathcal{S}$. We further refine the latent features $\boldsymbol{z}$ using transformers with self-attention.
To obtain shape information $\mathcal{S}$, we use point cloud $\mathcal{P}=\{\boldsymbol{p}_{i}\in\mathbb{R}^{3}\}$ and sparse view RGB images =$\mathcal{I}=\{\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}\}$. We encode the location of points into positional embeddings, resulting in features in $\mathbb{R}^{|\mathcal{P}|\times d}$. Meanwhile, a ViT network [17, 72] embeds each image into features of shape $\mathbb{R}^{h\times w\times d}$, resulting in the features of all images in $\mathbb{R}^{(|\boldsymbol{I}|\times h\times w)\times d}$. We use two transformers with cross-attention to aggregate information from them, respectively. We initialize the latent features $\bar{\boldsymbol{z}}\in\mathbb{R}^{n\times d}$ as the encoding of the points sampled using the farthest point sampling [114]. After the feature aggregation, we use an MLP to map the features to a lower-dimensional space $\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d_{0}}$, facilitating efficient latent diffusion.
Note that attention is invariant to the permutation of features. We also avoid pose-dependent operations [42, 43] in our design to enhance generality and extendability.
@@
The upsampling module is designed to increase both the length and feature dimension of the latent code, from $\boldsymbol{z}_{0}\in\mathbb{R}^{n\times d_{0}}$ to $\boldsymbol{z}_{l}\in\mathbb{R}^{M\times d}$. In detail, we first use a learnable query $\boldsymbol{y}$ to aggregate information from the latent code $\boldsymbol{z}_{0}$ using transformers with cross-attention. Then, we increase its channel dimension to $d$ using self-attention transformers, leading to the characteristic $\boldsymbol{z}_{1}\in\mathbb{R}^{n\times d}$. The architecture follows Perceiver-IO [39]. We then use an MLP to increase the feature dimension of $\boldsymbol{z}_{1}$ into $\frac{M}{n}d$, leading to latent features in $\mathbb{R}^{n\times\frac{M}{n}d}$. We reshape the features into $\boldsymbol{z}_{2}\in\mathbb{R}^{M\times d}$, which is a pixel shuffle process [85, 105]. We then use another set of transformers with self-attention to get the output latent features $\boldsymbol{z}_{l}$. This process is summarized as follows:
$\boldsymbol{z}_{1}=\text{SelfAttn}(\text{CrossAttn}(\boldsymbol{y},\boldsymbol{z}_{0})),\quad\boldsymbol{z}_{2}=\text{PixelShuffle}(\text{MLP}(\boldsymbol{z}_{1})),\quad\boldsymbol{z}_{l}=\text{SelfAttn}(\boldsymbol{z}_{2}).$ (6)
Again, we omit the feed-forward MLP in the transformers for simplicity. We then decode the patch centers using transformers with self-attention:
@@
Training. Similar to existing works [48, 105, 97, 121], we utilize a 3D dataset for supervision. We first regularize the patch center to adhere to the 3D surface geometry:
@@
where $\{\boldsymbol{\mathrm{\mu}}_{ij}\}_{i=1,j=1}^{M,S}$ represents the generated point cloud by sampling $S$ points in each of the $M$ patches. When computing the loss for supervision, we can generate point clouds with different resolutions by varying $S$, thanks to Atlas Gaussians’ ability to dynamically generate a variable number of 3D Gaussians.
While surface points are independent samples of the continuous geometry, 3D Gaussians are interdependent because one 3D Gaussian can affect the attributes of its neighbors during alpha blending. To ensure consistency and achieve deterministic results, during rendering we employ a uniform $\alpha\times\alpha$ grid sampling in the UV space instead of random sampling, resulting in $N=\alpha^{2}M$ 3D Gaussians. We use the differentiable renderer from [45] to render $V$ views of RGB, alpha, and depth images, which is supervised by mean square error:
@@
where $\hat{\boldsymbol{I}}_{\text{rgb}}$, $\hat{\boldsymbol{I}}_{\text{alpha}}$, and $\hat{\boldsymbol{I}}_{\text{depth}}$ are the predictions, $\boldsymbol{I}_{\text{rgb}}$, $\boldsymbol{I}_{\text{alpha}}$, and $\boldsymbol{I}_{\text{depth}}$ are the ground truths. To improve the visual fidelity, we also employ the LPIPS loss [117, 97] $\mathcal{L}_{\text{LPIPS}}(\hat{\boldsymbol{I}}_{\text{rgb}},\boldsymbol{I}_{\text{rgb}})$. The total loss function for the VAE is:
@@ 3.3 Stage 2: latent diffusion model
The proposed VAE provides a low-dimensional latent code that can be mapped to Atlas Gaussians. We employ the EDM [44] framework for latent diffusion. EDM aims to learn a denoising network $D_{\theta}(\boldsymbol{z};\sigma,\mathcal{C})$ to convert the Gaussian distribution to the empirical distribution $p_{\text{data}}$ defined by $\boldsymbol{z}$, where $\theta$ is the network parameters, $\sigma$ is the noise level sampled from a pre-defined distribution $p_{\text{train}}$, and $\mathcal{C}$ denotes the optional condition. $D_{\theta}(\boldsymbol{z};\sigma,\mathcal{C})$ is parameterized using a $\sigma$-dependent skip connection:
@@
where $\sigma\sim p_{\text{train}},\boldsymbol{z}\sim p_{\text{data}},\boldsymbol{n}\sim\mathcal{N}(0,\sigma^{2}I)$. Readers interested in details on $c_{\text{skip}}(\sigma)$, $c_{\text{out}}(\sigma)$, $c_{\text{in}}(\sigma)$, $c_{\text{noise}}(\sigma)$, and parameterization of the weight function $\lambda(\sigma)$ may refer to EDM [44].
Following [114], our denoising network $D_{\theta}(\boldsymbol{z};\sigma,\mathcal{C})$ consists of a series of blocks, each containing a self-attention layer followed by a cross-attention layer. Each block can be represented as:
where $\boldsymbol{z}^{(0)}=\boldsymbol{z}$ is the initial input to the network, $l$ is the number of blocks. In unconditional generation, we designate $\mathcal{C}$ as a learnable parameter. Note that more advanced architectures [74] could also be employed for the denoising network.
4 Experiments
We first introduce our experiment setup, and then show the results for both unconditional and conditional generation.
4.1 Experimental setup
Following most existing methods [6, 21, 69, 2, 8, 5, 48], we benchmark unconditional single-category 3D generation on ShapeNet [7]. We use the training split from SRN [89], which comprises 4612, 2151, and 3033 shapes in the categories Chair, Car, and Plane, respectively. We render 76 views for each shape using Blender [14] with the same intrinsic matrix as [5, 48]. Fréchet Inception Distance (FID@50K) and Kernel Inception Distance (KID@50K) are used for evaluation. In addition, we experiment with text-conditioned 3D generation on Objaverse [15]. We use the renderings from G-buffer Objaverse [77] and the captions from Cap3D [64]. Due to limited computational resources, we select a subset with around 6K 3D shapes. Please refer to Appendix A for implementation details.
4.2 Unconditional 3D generation
Table 1 presents the quantitative comparisons between our method and baseline approaches. Our method outperforms all baseline approaches, including EG3D [6], GET3D [21], DiffRF [69], RenderDiffusion [2], and SSDNeRF [8]. Our approach also outperforms concurrent work LN3Diff [48], which utilizes the triplane representation for 3D shapes. Our results demonstrate significant improvements over LN3Diff, particularly in the ShapeNet Chair category, which features greater geometric, structural, and textural complexity. In the other two categories, our method achieves better average performance, further highlighting the robustness and efficacy of our approach. We also include the qualitative comparison in Figure 5.

### v3

@@ 3.2 Stage 1: VAE
We design a VAE to link the latent space and the 3D space. The overall pipeline is shown in Figure 3.
Encoder. Following the latent set representation in Zhang et al. (2023a), the encoder takes shape information as input and outputs a latent set $\boldsymbol{z}_{0}\in\mathbb{R}^{n\times d_{0}}$, where $n$ is the size of the set in the latent set representation and $d_{0}$ is the latent dimension. The shape information includes point cloud $\mathcal{P}=\{\boldsymbol{p}_{i}\in\mathbb{R}^{3}\}$ and sparse view RGB images =$\mathcal{I}=\{\boldsymbol{I}_{i}\in\mathbb{R}^{H\times W\times 3}\}$. We encode the location of points into positional embeddings, resulting in features in $\mathbb{R}^{|\mathcal{P}|\times d}$. Meanwhile, a ViT network Dosovitskiy et al. (2021); Oquab et al. (2024) embeds each image into features of shape $\mathbb{R}^{h\times w\times d}$, resulting in the features of all images in $\mathbb{R}^{(|\boldsymbol{I}|\times h\times w)\times d}$. We then initialize the latent features $\bar{\boldsymbol{z}}\in\mathbb{R}^{n\times d}$ as the encoding of the points sampled using the farthest point sampling Zhang et al. (2023a) and use transformers with cross-attention to aggregate input shape information. After feature aggregation, we use an MLP to map the features to a lower-dimensional space $\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d_{0}}$, facilitating efficient latent diffusion. The process is summarized as follows:
$\boldsymbol{z}^{\prime}=\text{CrossAttn}(\bar{\boldsymbol{z}},\mathcal{P}),\quad\boldsymbol{z}^{\prime\prime}=\text{CrossAttn}(\boldsymbol{z}^{\prime},\mathcal{I}),\quad\boldsymbol{z}^{\prime\prime\prime}=\text{SelfAttn}(\boldsymbol{z}^{\prime\prime}),\quad\boldsymbol{z}_{0}=\text{MLP}(\boldsymbol{z}^{\prime\prime\prime}).$ (6)
In this notation, we omit the feed forward MLP of the transformers for simplicity. Intuitively, the query $\bar{\boldsymbol{z}}$ is updated by aggregating the input point features and image features.
Note that attention is invariant to the permutation of features. We also avoid pose-dependent operations Jiang et al. (2024); Jun & Nichol (2023) in our design to enhance generality and extendability.
@@
The upsampling module is designed to increase both the length and the feature dimension of the latent code, from $\boldsymbol{z}_{0}\in\mathbb{R}^{n\times d_{0}}$ to $\boldsymbol{z}_{l}\in\mathbb{R}^{M\times d}$. In detail, we first use a learnable query $\boldsymbol{y}$ to aggregate information from the latent code $\boldsymbol{z}_{0}$ using transformers with cross-attention. Then, we increase its channel dimension to $d$ using self-attention transformers, leading to the characteristic $\boldsymbol{z}_{1}\in\mathbb{R}^{n\times d}$. We then use an MLP to increase the feature dimension of $\boldsymbol{z}_{1}$ into $\frac{M}{n}d$, leading to latent features in $\mathbb{R}^{n\times\frac{M}{n}d}$. We reshape the features into $\boldsymbol{z}_{2}\in\mathbb{R}^{M\times d}$, which is a pixel shuffle process Shi et al. (2016); Xu et al. (2024a). We then use another set of transformers with self-attention to obtain the output latent features $\boldsymbol{z}_{l}$. This process is summarized as follows:
$\boldsymbol{z}_{1}=\text{SelfAttn}(\text{CrossAttn}(\boldsymbol{y},\boldsymbol{z}_{0})),\quad\boldsymbol{z}_{2}=\text{PixelShuffle}(\text{MLP}(\boldsymbol{z}_{1})),\quad\boldsymbol{z}_{l}=\text{SelfAttn}(\boldsymbol{z}_{2}),$ (7)
where $\boldsymbol{y}$ is initialized using a Gaussian distribution. Again, we omit the feed-forward MLP in the transformers for simplicity. We then decode the patch centers using transformers with self-attention:
@@
Training. Similar to existing methods Lan et al. (2024); Xu et al. (2024a); Zou et al. (2023), we utilize a 3D dataset for supervision. We first regularize the patch center to adhere to the 3D surface geometry:
@@
where $\{\boldsymbol{\mathrm{\mu}}_{ij}\}_{i=1,j=1}^{M,S}$ represents the generated point cloud by sampling $S$ points in each of the $M$ patches. When computing the loss for supervision, we can generate point clouds with different resolutions by varying $S$, thanks to Atlas Gaussians’ ability to dynamically generate a variable number of 3D Gaussians. Note that patches may overlap, similar to AtlasNet. $\mathcal{L}_{\text{center}}$ and $\mathcal{L}_{\boldsymbol{\mathrm{\mu}}}(S)$ encourage the patches to distribute more uniformly across the surface, thereby making better use of the 3D Gaussians.
While surface points are independent samples of the continuous geometry, 3D Gaussians are interdependent because one 3D Gaussian can affect the attributes of its neighbors during alpha blending. To ensure consistency and achieve deterministic results, during rendering we employ a uniform $\alpha\times\alpha$ grid sampling in the UV space instead of random sampling, resulting in $N=\alpha^{2}M$ 3D Gaussians. We use the differentiable renderer from Kerbl et al. (2023) to render $V$ views of RGB, alpha, and depth images, which is supervised by mean square error:
@@
where $\hat{\boldsymbol{I}}_{\text{rgb}}$, $\hat{\boldsymbol{I}}_{\text{alpha}}$, and $\hat{\boldsymbol{I}}_{\text{depth}}$ are the predictions, $\boldsymbol{I}_{\text{rgb}}$, $\boldsymbol{I}_{\text{alpha}}$, and $\boldsymbol{I}_{\text{depth}}$ are the ground truths. To improve the visual fidelity, we also employ the LPIPS loss Zhang et al. (2018); Tang et al. (2024) $\mathcal{L}_{\text{LPIPS}}(\hat{\boldsymbol{I}}_{\text{rgb}},\boldsymbol{I}_{\text{rgb}})$. The total loss function for the VAE is:
@@ 3.3 Stage 2: latent diffusion model
The proposed VAE provides a low-dimensional latent code that can be mapped to Atlas Gaussians. We employ EDM Karras et al. (2022) for latent diffusion. Our denoising network follows the same architecture as Zhang et al. (2023a), which consists of a series of transformer blocks:
where $\boldsymbol{z}^{(0)}=\boldsymbol{z}_{0}$ is the initial input to the network, $l$ is the number of blocks. In unconditional generation, we designate $\mathcal{C}$ as a learnable parameter. For text-conditioned generation, $\mathcal{C}$ is set as the CLIP embedding Radford et al. (2021) of the input text prompts. Since the LDM network design is not our main contribution, we provide the implementation details in Appendix A.3.
4 Experiments
We first introduce our experimental setup, and then show the results for both unconditional and conditional generation.
4.1 Experimental setup
Following most existing methods Gao et al. (2022); Müller et al. (2023); Chen et al. (2023a); Lan et al. (2024), we benchmark unconditional single-category 3D generation on ShapeNet Chang et al. (2015). We use the training split from SRN Sitzmann et al. (2019), which comprises 4612, 2151, and 3033 shapes in the categories Chair, Car, and Plane, respectively. We render 76 views for each shape using Blender Community (2018) with the same intrinsic matrix as Cao et al. (2024); Lan et al. (2024). Fréchet Inception Distance (FID@50K) and Kernel Inception Distance (KID@50K) are used for evaluation. In addition, we experiment with text-conditioned 3D generation on Objaverse Deitke et al. (2022). We use the renderings from G-buffer Objaverse Qiu et al. (2023) and the captions from Cap3D Luo et al. (2023). Due to limited computational resources, we select a high-quality subset with around 18K 3D shapes. We use CLIP score Radford et al. (2021); Hessel et al. (2021), FID and KID for evaluation. Please refer to Appendix A.1 for more implementation details.
4.2 Unconditional 3D generation
Table 1 presents the quantitative comparison between our method and baseline approaches. Our method outperforms all baseline approaches, including EG3D Chan et al. (2022), GET3D Gao et al. (2022), DiffRF Müller et al. (2023), RenderDiffusion Anciukevičius et al. (2023), SSDNeRF Chen et al. (2023a) and LN3Diff Lan et al. (2024). We also include the qualitative comparison with LN3Diff in Figure 5. Our results demonstrate significant improvements over LN3Diff, particularly in the ShapeNet Chair category, which features greater geometric, structural, and textural complexity. The results highlight the robustness and efficacy of our approach.
[…]
Figure 5: The comparison between LN3Diff and our method on three ShapeNet categories.
Table 1: Evaluation of single-category unconditional generation on ShapeNet.
Method Chair Car Plane FID@50K KID@50K(%) FID@50K KID@50K(%) FID@50K KID@50K(%) EG3D Chan et al. (2022) 26.09 1.1 33.33 1.4 14.47 0.5 GET3D Gao et al. (2022) 35.33 1.5 41.41 1.8 26.80 1.7 DiffRF Müller et al. (2023) 99.37 4.9 75.09 5.1 101.79 6.5 RenderDiffusion Anciukevičius et al. (2023) 53.30 6.4 46.50 4.1 43.50 5.9 SSDNeRF Chen et al. (2023a) 65.04 3.0 47.72 2.8 21.01 1.0 LN3Diff Lan et al. (2024) 16.90 0.47 17.60 0.49 8.84 0.36 Ours 9.90 0.35 12.15 0.45 8.09 0.21
4.3 Conditional 3D generation
We evaluate our method and baseline approaches He et al. (2024); Lan et al. (2024); Tang et al. (2024); Jun & Nichol (2023) on text-conditioned 3D generation using Objaverse. The qualitative results are presented in 6, where all text prompts are sourced from the original baseline papers. As shown in Figure 6, the image-based generalizable 3D reconstruction method Tang et al. (2024) sometimes produces shapes with significant artifacts, primarily due to its reliance on a multi-view generation network, which frequently suffers from multi-view inconsistency. In contrast, our method learns directly from 3D data, enabling the generation of consistent novel views. Additionally, our method produces higher-quality 3D shapes compared to Shap-E and LN3Diff, both of which use the VAE + LDM paradigm. However, these methods depend on volume rendering during training, which is typically restricted to low resolution. Our approach instead leverages the more efficient 3D Gaussian representation. Compared to GVGEN, our generated shapes exhibit finer details, owing to the proposed Atlas Gaussians representation, which is capable of decoding a significantly larger number of 3D Gaussians. Quantitative results are provided in Table 2, where our method achieves the best performance across all metrics and has the shortest inference time. This improvement is due to our efficient decoder design, which effectively links the low-dimensional latent space with 3D space. Additional results can be found in Appendix B.
[…]
Figure 6: Comparison of text-conditioned 3D generation on Objaverse. From left to right: GVGEN He et al. (2024), LN3Diff Lan et al. (2024), LGM Tang et al. (2024), Shap-E Jun & Nichol (2023), and our method. All text prompts are sourced from the original baseline papers.
[…]
Figure 7: (Left) Our generated results demonstrate significant diversity. (Right) Our generated results align closely with the text prompts, allowing for strong controllability. In the second row of each group, we present the nearest neighbors (NN) from the training dataset.
In Figure 7, we provide a detailed analysis of our method. In Figure 7 (Left), we present text-conditioned generation results with different random seeds, demonstrating that our method produces highly diverse outputs. In Figure 7 (Right), we show that our model can be robustly controlled using different text prompts. The generated results are also significantly different from their nearest neighbors in the training dataset, highlighting the model’s ability to generate novel content.
Table 2: Evaluation of text-conditioned 3D generation on Objaverse.
[…]
Atlas Gaussians (Full) 26.56 (a) linear weights 26.13 (b) no disentangle 25.77 (c) no global feature 26.18 (d) decode $\boldsymbol{\mathrm{s}}$,$\boldsymbol{\mathrm{r}}$,${\mathrm{o}}$ from $\boldsymbol{f}_{i}$ 26.55
4.4 Ablation study
Number of 3D Gaussians and number of patches. In Table 4, we analyze the effects of both the number of patches and the total number of 3D Gaussians on LPIPS and memory usage (MEM, which represents the additional GPU memory required for each unit increase in batch size). The results indicate that increasing either the number of patches or the number of 3D Gaussians improves LPIPS. When the number of 3D Gaussians is fixed (32k or 100k), increasing the number of patches from 2,048 to 4,096 results in a 0.1-0.2 improvement in LPIPS. However, when the number of patches is fixed (2,048 or 4,096), increasing the number of 3D Gaussians from 32k to 100k leads to a more significant improvement in LPIPS while requiring less additional GPU memory. Importantly, the network parameters remain unchanged despite the increase in 3D Gaussians. This comparison demonstrates that increasing the number of 3D Gaussians through Atlas Gaussians decoding is more efficient and effective than increasing the number of patches. Notably, the Atlas Gaussians representation combines explicit patch generation with implicit interpolation to generate 3D Gaussians, offering a balanced trade-off between compact implicit and fast explicit representations.
Disentangle geometry and texture features. Atlas Gaussians use separate branches to learn geometry and appearance features. To validate the effectiveness of this design, we experimented with an alternative network version where a single set of features is shared for both geometry and appearance. As shown in Table 4, performance decreases considerably with this shared feature approach. This shows that our design facilitates more effective learning. Note that both the geometry and appearance features are generated from the shared latent $\boldsymbol{z}_{0}$, and disentanglement is performed during feature generation rather than in the latent space.
Parameter decoding scheme. In Gaussian Splatting, opacity, scale, and rotation also influence the geometry. We experimented with decoding opacity, scale, and rotation from the geometry features and found that the performance is nearly identical to when these attributes are decoded from the appearance features. This result indicates that the key factor is decoding the Gaussian centers and RGB colors using distinct, separate branches.
5 Conclusion
Acknowledgment. We thank Yushi Lan for his invaluable assistance in the evaluation. Qixing Huang acknowledges the support from NSF IIS-2047677, IIS-2413161, and GIFTs from Adobe and Google.
Appendix A Implementation details
A.1 Additional details of the dataset
[…]
We randomly selected 250 text prompts for evaluation, ensuring that each testing prompt differs from the training data. For each object, we rendered 24 views uniformly, following the G-buffer Objaverse protocol, resulting in a total of 6K images. FID and KID are computed based on the 2D image feature space.
A.2 Additional details of VAE
The VAE training consists of two stages. In the first stage, $\lambda_{r}$ in Eq. 12 is set to 0. Rendering occurs only in the second stage with $\lambda_{r}$ set to 1. In both stages, $\lambda_{KL}$ maintains $1e^{-4}$. As shown in Eq. 12, all other loss weights are set to 1 for simplicity. Our method is robust to hyperparameters; doubling or halving $\lambda_{r}$ in Eq. 12 results in almost identical loss curves.
We utilize a sparse point cloud of 2048 points and 4 input views, each of size $224\times 224$. $M=2048$ patches are used in all experiments. For the latent, $n=512$, $d=512$. We set $d_{0}$ to 4 for ShapeNet and $d_{0}$ to 16 in Objaverse. In ShapeNet, $\alpha$ is set to 4, resulting in $N=32768$ 3D Gaussians. In Objaverse, $\alpha=7$. Due to the real-time rendering capabilities of the 3D Gaussians, we are able to produce an output image of size $512\times 512$. We use the EMD implementation from Liu et al. (2019), which supports fast and efficient computation even with up to 8192 points. All networks are trained on 8 Tesla V100 GPUs for 1000 epochs using the AdamW optimizer Loshchilov & Hutter (2019) with the one-cycle policy. Our VAE is trained using mixed precision (fp16) and supports a batch size of 8 per GPU. For instance, when trained on the ShapeNet dataset, it requires approximately 22GB of memory per GPU, making it accessible to a wide range of laboratories. Notably, the model does not require the latest A100 GPUs for training. Furthermore, the training process is completed in less than 30 hours. For Objaverse, the VAE training takes about 6 days.
In terms of network parameters, our model increases the number of Gaussians without adding extra network parameters. Specifically, for Objaverse, our model requires only 142M network parameters to generate 100K Gaussians, while LGM Tang et al. (2024) requires 415M parameters for generating 64K Gaussians. This demonstrates that our representation is more efficient than purely explicit representations like LGM, offering a nice trade-off between compact implicit and fast explicit representations.
A.3 Additional details of LDM
@@
We adopt the EDM Karras et al. (2022) framework for latent diffusion. EDM aims to learn a denoising network $D_{\theta}(\boldsymbol{z};\sigma,\mathcal{C})$ to convert the Gaussian distribution to the empirical distribution $p_{\text{data}}$ defined by $\boldsymbol{z}$, where $\theta$ is the network parameters, $\sigma$ is the noise level sampled from a predefined distribution $p_{\text{train}}$, and $\mathcal{C}$ denotes the optional condition. $D_{\theta}(\boldsymbol{z};\sigma,\mathcal{C})$ is parameterized using a $\sigma$-dependent skip connection:
@@
where $\sigma\sim p_{\text{train}},\boldsymbol{z}\sim p_{\text{data}},\boldsymbol{n}\sim\mathcal{N}(0,\sigma^{2}I)$. Readers interested in details on $c_{\text{skip}}(\sigma)$, $c_{\text{out}}(\sigma)$, $c_{\text{in}}(\sigma)$, $c_{\text{noise}}(\sigma)$, and parameterization of the weight function $\lambda(\sigma)$ may refer to EDM Karras et al. (2022).
For the latent diffusion model, we set $l=12$ for ShapeNet and $l=24$ for Objaverse. The final latents are obtained via 40 denoising steps. For text-conditioned 3D generation, we use CLIP to encode the input text prompts. In addition, we adopt classifier-free guidance. We randomly drop the conditioning signal with a probability of $10\%$ and set the guidance scale to 3.5 during sampling. Note that more advanced architectures Peebles & Xie (2022) could also be employed for the denoising network.
Appendix B Additional results
In Figure 8, we present additional results of text-conditioned 3D generation using our method.
Figure 9 shows the results of real-image-conditioned 3D generation. In this experiment, we simply replace the CLIP text encoder with the CLIP image encoder and train the latent diffusion model using the same dataset. Given a real image, we first apply an off-the-shelf tool to remove the background and then use the processed image as the conditional input for 3D generation. The results demonstrate that our method can still produce reasonable outcomes.